OpenStudy (anonymous):

How do I complete this problem? 7√3 - 4√6 + √48 - √54 √48 = 4√3 √54 = 3√6

OpenStudy (anonymous):

Ok the idea is to simplify the \[\sqrt{48} \space and \space \sqrt{54}\]

OpenStudy (anonymous):

7√3 + 4√3 - 4√6 - 3√6 = 11√3 - 7√6

OpenStudy (anonymous):

So sqrt 48 can be written like this \[\sqrt{16 * 3} = \sqrt{16}*\sqrt{3}=4\sqrt{3}\]

OpenStudy (anonymous):

Now to simplify the sqrt(54).. 54 can be rewritten as\[\sqrt{9*6} = \sqrt{9}*\sqrt{6} = 3\sqrt{6}\] so now you have \[7\sqrt{3} - 4\sqrt{6} + 4\sqrt{3} - 3\sqrt{6}\]

OpenStudy (anonymous):

Which would be \[11\sqrt{3} - 7\sqrt{6}\]